find an equation through each given pair of points. Write the answer in standard from using only integers. (-1,1) (1,-1)
Answer by stanbon(75887)

find an equation through each given pair of points. Write the answer in standard from using only integers. (-1,1) (1,-1)
---------------------
slope = (-1-1)/(1--1) = -2/2 = -1
-------------------------------------
Form: y = mx + b
Using m = -1 and point (-1,1), solve for "b":
1 = -1*-1 + b
b = 0
----
Equation: y = -x
=======================
Cheers,
Stan H.